OpenStudy (anonymous):

What is the simplified form of x^2+6x+8 over 4-x^2

OpenStudy (theviper):

Click on Bump it will go up :)

OpenStudy (anonymous):

\[\huge x^2+6x+8 \over \huge4-x^2\]

OpenStudy (callisto):

Way to factor the two expressions: x^2+6x+8 = x^2 + 2x + 4x + 8 = x(x+2) + 4(x+2) = (x+4)(x+2) 4-x^2 =2^2 - x^2 Use the identity a^b - b^2 = (a+b)(a-b) = (2-x)(2+x)